OpenStudy (anonymous):

Factored form of (36x^2-24x)(12x-18)

OpenStudy (anonymous):

(72x)(3x-2)(2x-3)

OpenStudy (anonymous):

whenever you factor you want to try and pull out the highest value from each term

OpenStudy (anonymous):

belangerk....factor each term, so 36x^2 - 24x = 12x(3x - 2) and 12x - 18 = 6(2x - 3) So we pulled out a 12x and a 6, when multiplied, that gives you 72x. So your final answer is 72x(3x - 2)(2x - 3)